Bob and Mia both like making bread and jam. In one day Bob can make 10 loafs of bread OR
20 jars of jam. In one day Mia can make 6 loafs of bread OR 36 jars of jam.

- If Bob only makes jam and Mia only makes bread, how many loafs of bread and jars of jam would they
have between them?
- Assume Bob and Mia specialize in their comparative advantage good, and then agree to trade 3 loafs
of bread for 12 jars of jam. How many loafs of bread and jars of jam do they each have now?
- Has specializing in their comparative advantage good made them better off overall? Explain your
answer fully, including explaining what is meant by ‘comparative advantage’.
Answer #1

Each day production table:

Bread Jam
Bob 10 20
Mia 6 36

In one day Bob can make 10 loafs of bread OR 20 jars of jam. In one day Mia can make 6 loafs of bread OR 36 jars of jam.

If Bob only makes jam and Mia only makes bread, then 20 jams and 6 bread they both can make together.

--------

Opportunity cost of producing Bread in terms of Jam Opportunity cost of producing jam in terms of bread
Bob (20/10) = 2 jam per bread (10/20) = 0.5 bread per jam
Mia (36/6) = 6 jam per bread (6/36) = 0.167 bread per jam

Bob opportunity cost of producing bread in terms of jam is lower than Mia's opportunity cost of prodcuing bread in term of jam. It means Bob has comparative advantage in production of bread and should specialize in bread production.

In similar fashion, Mia has comparative advantage in prodution of jam and should specialize in Jam production.

After specilization, Bob will produce only Bread and Mia will produce only Jam.

Bob can make 10 loafs to bread per day and Mia can make 36 jars of jam per day.

They agree to trade 3 loafs of bread for 12 jars of jam.

After trade, Bob will have 7 loafs of bread (i.e., 10 - 3 = 7) and 12 jars of jam.

and, Mia will have 3 loafs of bread and 24 jars of jam (i.e., 36 -12 =24)

---------------

Yes, specializing in their comparative advantage good made them better off overall.

Before specialization they both were making 6 loafs of bread and 20 jars of jam.

After specialization and trade, they both have 10 loafs of bread (i.e., 7 + 3=10) and 36 jars of jam (i.e., 12 + 24 =36)

A person is said to be have a comparative advanatge in a good when the opportunity cost of produicng that good in terms of other good is less than the other person opportunity cost in same good.